OpenStudy (mackenzie2013):

The ratio of shirts to pants is 3 to 2. If there are 6 shirts, how many pants are there?

OpenStudy (anonymous):

NO WAIT

OpenStudy (anonymous):

4

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

should i explain?

OpenStudy (anonymous):

\[S/P=3/2 \implies P=S\times(2/3)=6*(2/3)=4\]

OpenStudy (anonymous):

duh

OpenStudy (comm.dan):

Hey @Mackenzie2013 do you know how to start solving the problem?

OpenStudy (anonymous):

You have to explain it's in the policy lol

OpenStudy (anonymous):

3/2 = 6/x x=6x2/3 x=4

OpenStudy (mackenzie2013):

I do !Thanks for the answers and explaining ! Was checking to make sure it was right!

OpenStudy (anonymous):

little talks of monsters and men :3 i did

OpenStudy (anonymous):

That song <3

OpenStudy (anonymous):

its nice ^^ love the video